# YITP-25-79, IPMU25-0028, RIKEN-iTHEMS-Report-25

### v1

@@ 4.3.2 $\eta_{1}$ correction
where the coefficients $\alpha_{j}^{-}$ are proportional to $\eta_{1}$. Here, we do not show the full expression of $\alpha_{j}^{-}$, but the formula for $\ell=2$ can be found in [34]. [^17] Note that $\delta V_{\ell}^{-}$, $r_{\rm H}$, and $\epsilon_{1}$ therein should be understood, respectively as $r_{g}^{-2}\delta V_{\ell}^{-}$, $r_{g}$, and $-\eta_{1}$ in our notation. Following the same calculation as before, one can straightforwardly obtain the corrections to the TDN for $\ell=2$ as
@@ 5 Conclusions
There are several future directions that we would like to explore. First, an obvious extension of this work is to generalize our parametrized formalism to a stationary and axisymmetric BH background. [^19] In fact, this can be straightforwardly generalized to a rotating background since the MST solution of the Teukolsky equation was found in [59, 57, 69]. Therefore, with that solution, one can, in principle, formulate the parametrized formalism for the modified Teukolsky equation. By doing so, it would generally capture the effects of BH spin in both the static TLNs and the TDNs. Second, it is interesting to extend our formalism to systems where there are couplings among different physical degrees of freedom. This study would reveal a rich structure of the tidal response in the alternative theories of gravity. Third, it would be nice to formulate the parametrized formalism taking into account the higher-order corrections in frequency in the tidal response function. This would require the matching with the solutions obtained from the worldline EFT. [^20] Indeed, clarifying the connection with the point-particle EFT would also resolve the relation between TLNs and Green’s functions in BH perturbations [93]. We leave this investigation to our future work [94]. Finally, with the increasing number of future GW observations, it is important to study the impact of tidal dissipation on the waveform, as this could provide a way to constrain deviations from GR in the vicinity of BH. These directions would pave the way for a deeper understanding of tidal responses in alternative theories of gravity and their potential observational signatures.
@@ Acknowledgements
V.Y. is grateful for the hospitality of the external program APCTP-GW2025 held at Academia Sinica, Taipei, Taiwan, where part of this work was carried out. This work was supported in part by World Premier International Research Center Initiative (WPI), MEXT, Japan. The work of H.K. was supported by JST (Japan Science and Technology Agency) SPRING, Grant No. JPMJSP2110. The work of S.M. was supported in part by JSPS (Japan Society for the Promotion of Science) KAKENHI Grant No. JP24K07017. The work of N.O. was supported by JSPS KAKENHI Grant No. JP23K13111 and by the Hakubi project at Kyoto University. The work of K.T. was supported by JSPS KAKENHI Grant No. JP23K13101. The work of V.Y. is supported by grants for development of new faculty staff, Ratchadaphiseksomphot Fund, Chulalongkorn University.
@@ A.1 Low-frequency expansion of the MST coefficients $a_{n}^{\nu}$
$1$ $-i\varepsilon\frac{(\ell+3)^{2}}{2(\ell+1)(2\ell+1)}+\varepsilon^{2}\frac{(\ell+3)^{2}}{2(\ell+1)^{2}(2\ell+1)}+{\cal O}(\varepsilon^{3})$
$-1$ $-i\varepsilon\frac{(\ell-2)^{2}}{2\ell(2\ell+1)}-\varepsilon^{2}\frac{(\ell-2)^{2}}{2\ell^{2}(2\ell+1)}+{\cal O}(\varepsilon^{3})$
$2$ $-\varepsilon^{2}\frac{(\ell+3)^{2}(\ell+4)^{2}}{4(\ell+1)(2\ell+1)(2\ell+3)^{2}}+{\cal O}(\varepsilon^{3})$
$-2$ $-\varepsilon^{2}\frac{(\ell-3)^{2}(\ell-2)^{2}}{4\ell(2\ell-1)^{2}(2\ell+1)}+{\cal O}(\varepsilon^{3})$
@@ A.3 MST solutions for $\ell=2$
$\delta{\cal F}_{\nu}^{-,\,{\rm Bare}}(\varepsilon)$ $=\mathopen{}\mathclose{{}\left(\delta C_{+}^{-}+\delta C_{-}^{-}}\right)\times\frac{\mathopen{}\mathclose{{}\left(\text{the coefficient of $x^{-\nu}$ in $\psi_{-\nu-1}^{(0)}(x)$}}\right)}{\mathopen{}\mathclose{{}\left(\text{the coefficient of $x^{\nu+1}$ in $\psi_{\nu}^{(0)}(x)$}}\right)}$
@@ Appendix B Logarithmic running of tidal Love numbers
[…]
@@ C.2 Tidal response of stealth Schwarzschild background
${\cal F}_{\ell}^{-,\,{\rm Ren}}=ir_{g}\tilde{\omega}{\cal N}_{\ell,{\rm GR}}+(r_{g}\tilde{\omega})^{2}{\cal K}^{(2)}_{\ell,{\rm GR}}+\cdots\;,$ (C.10)
@@
Hence, when $\alpha_{T}<0$ and $r_{g}>r_{s}$ (subluminal propagation), we have ${\cal N}_{\ell}^{-}>{\cal N}_{\ell,{\rm GR}}$ and $\big{|}{\cal K}_{\ell}^{(2)-}\big{|}>\big{|}{\cal K}^{(2)}_{\ell,{\rm GR}}\big{|}$. On the other hand, when $\alpha_{T}>0$ and $r_{g}<r_{s}$ (superluminal propagation), ${\cal N}_{\ell}^{-}$ and $\big{|}{\cal K}_{\ell}^{(2)-}\big{|}$ are smaller than their corresponding GR values.

### v2

@@ 4.3.2 $\eta_{1}$ correction
where the coefficients $\alpha_{j}^{-}$ are proportional to $\eta_{1}$. Here, we do not show the full expression of $\alpha_{j}^{-}$, but the formula for $\ell=2$ can be found in [22]. [^17] Note that $\delta V_{\ell}^{-}$, $r_{\rm H}$, and $\epsilon_{1}$ therein should be understood, respectively as $r_{g}^{-2}\delta V_{\ell}^{-}$, $r_{g}$, and $\eta_{1}$ in our notation. Following the same calculation as before, one can straightforwardly obtain the corrections to the TDN for $\ell=2$ as
@@ 5 Conclusions
There are several future directions that we would like to explore. First, an obvious extension of this work is to generalize our parametrized formalism to a stationary and axisymmetric BH background. [^19] In fact, this can be straightforwardly generalized to a rotating background since the MST solution of the Teukolsky equation was found in [71, 69, 84]. Therefore, with that solution, one can, in principle, formulate the parametrized formalism for the modified Teukolsky equation (see Ref. [25] for the corrected TLNs of non-rotating BHs in this context). By doing so, it would generally capture the effects of BH spin in both the static TLNs and the TDNs. Second, it is interesting to extend our formalism to systems where there are couplings among different physical degrees of freedom. This study would reveal a rich structure of the tidal response in the alternative theories of gravity. Third, it would be nice to formulate the parametrized formalism taking into account the higher-order corrections in frequency in the tidal response function. This would require the matching with the solutions obtained from the worldline EFT. [^20] Indeed, clarifying the connection with the point-particle EFT would also resolve the relation between TLNs and Green’s functions in BH perturbations [105]. We leave this investigation to our future work [106]. Finally, with the increasing number of future GW observations, it is important to study the impact of tidal dissipation on the waveform, as this could provide a way to constrain deviations from GR in the vicinity of BH. These directions would pave the way for a deeper understanding of tidal responses in alternative theories of gravity and their potential observational signatures.
@@ Acknowledgements
V.Y. is grateful for the hospitality of the external program APCTP-GW2025 held at Academia Sinica, Taipei, Taiwan, where part of this work was carried out. This work was supported in part by World Premier International Research Center Initiative (WPI), MEXT, Japan. The work of H.K. was supported by JST (Japan Science and Technology Agency) SPRING, Grant No. JPMJSP2110. The work of S.M. was supported in part by JSPS (Japan Society for the Promotion of Science) KAKENHI Grant No. JP24K07017. The work of N.O. was supported by JSPS KAKENHI Grant No. JP23K13111 and by the Hakubi project at Kyoto University. The work of K.T. was supported by JSPS KAKENHI Grant No. JP23K13101. The work of V.Y. is supported by grants for development of new faculty staff, Ratchadaphiseksomphot Fund, Chulalongkorn University and by the National Science, Research and Innovation Fund (NSRF) via the Program Management Unit for Human Resources & Institutional Development, Research and Innovation Grant No. B39G680009.
@@ A.1 Low-frequency expansion of the MST coefficients $a_{n}^{\nu}$
11 $-i\varepsilon\frac{(\ell+3)^{2}}{2(\ell+1)(2\ell+1)}+\varepsilon^{2}\frac{(\ell+3)^{2}}{2(\ell+1)^{2}(2\ell+1)}+{\cal O}(\varepsilon^{3})$
1-1 $-i\varepsilon\frac{(\ell-2)^{2}}{2\ell(2\ell+1)}-\varepsilon^{2}\frac{(\ell-2)^{2}}{2\ell^{2}(2\ell+1)}+{\cal O}(\varepsilon^{3})$
22 $-\varepsilon^{2}\frac{(\ell+3)^{2}(\ell+4)^{2}}{4(\ell+1)(2\ell+1)(2\ell+3)^{2}}+{\cal O}(\varepsilon^{3})$
2-2 $-\varepsilon^{2}\frac{(\ell-3)^{2}(\ell-2)^{2}}{4\ell(2\ell-1)^{2}(2\ell+1)}+{\cal O}(\varepsilon^{3})$
@@ A.3 MST solutions for $\ell=2$
[…]
@@ Appendix B Logarithmic running of tidal Love numbers
Let us recall that in this paper we use the analytic continuation of $\ell$ to non-integer values ($\ell\to\nu=\ell+\Delta\ell$) in order to distinguish between the source and response terms. By doing so, the solution may develop a pole in the limit $\Delta\ell\to 0$. For the MST solution discussed in Sections 2, such a singularity appears in $\psi_{\nu+1}^{(0)}(x)$ as $C_{\ell}^{\rm div}(\varepsilon)x^{-\ell+\Delta\ell}/(2\Delta\ell)$ and in $\psi_{-\nu}^{(0)}(x)$ as $C_{\ell}^{\rm div}(\varepsilon)x^{-\ell-\Delta\ell}/(2\Delta\ell)$, where the coefficient $C_{\ell}^{\rm div}$ starts at the second order in $\varepsilon$, i.e., $C_{\ell}^{\rm div}(\varepsilon)\sim{\cal O}(\varepsilon^{2})$. If we let $\Delta\ell=\ell_{2}\varepsilon^{2}+{\cal O}(\varepsilon^{4})$, this would lead to a pole in the limit $\ell_{2}\to 0$. [^21] One way to interpret this pole is through the worldline EFT framework, where it arises from divergences in loop diagrams and their corresponding counterterms when using dimensional regularization. Notice that the factor $\ell_{2}$ does not appear in the denominator of the total horizon-ingoing solution. In the case of vacuum 4d GR, the bare static TLNs are identically zero, so that the counterterm vanishes, i.e., $C_{\ell}^{\rm div}\to 0$ as $\varepsilon\to 0$. This is indeed a consequence of the protection by hidden ladder symmetry (see, e.g., [108, 109, 110, 111, 112, 113]).
@@ C.2 Tidal response of stealth Schwarzschild background
${\cal F}_{\ell}^{-,\,{\rm Ren}}=ir_{g}\tilde{\omega}{\cal N}_{\ell,{\rm GR}}+(r_{g}\tilde{\omega})^{2}{\cal K}^{(2)}_{\ell,{\rm GR}}$$\cdots\;$, (C.10)
@@
[…]